OpenStudy (anonymous):

find any vertical asymptotes of the function limit of 1 divided by the quantity x minus 8 as x approaches 8 from the left

OpenStudy (anonymous):

\[f(x)=\frac{1}{x-8}\]?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

you know what it looks like?

OpenStudy (anonymous):

not exactly

OpenStudy (anonymous):

has a vertical asymptote at \(x=8\) for sure

OpenStudy (anonymous):

did you graph these in some pre-calc class?

OpenStudy (anonymous):

this is for my final assignment for my online class

OpenStudy (anonymous):

actually we can cheat easily, do it without knowing the graph

OpenStudy (anonymous):

oh well if it is on line lets cheat like crazy

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1%2F%28x-8%29

OpenStudy (anonymous):

you and i would become good friends lol

OpenStudy (anonymous):

you can see that as \(x\to 8^-\) it goes to \(-\infty\)

OpenStudy (anonymous):

you can see that from the picture right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok my motto "cheating is learning"

OpenStudy (anonymous):

awesome. that is now my new motto

OpenStudy (anonymous):

all rights reserved

OpenStudy (anonymous):

wait so is x=8 the only one?

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

only place where the denominator would be zero

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

and is it ok if you can help me with one last question

OpenStudy (anonymous):

ok one last question

OpenStudy (anonymous):

Find the fifth roots of 243(cos 240° + i sin 240°)

OpenStudy (anonymous):

lord i thought it as going to be a short one

OpenStudy (anonymous):

ok first the fifth root of 243 is 3 i think

OpenStudy (anonymous):

lol sorry.

OpenStudy (anonymous):

then divide \(240\) by \(5\)

OpenStudy (anonymous):

i get 48

OpenStudy (anonymous):

that would be 48

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so one answer is \[ 3(\cos (48°) + i \sin (48°))\]

OpenStudy (anonymous):

next we add 360 to 240 and divide by 5 again

OpenStudy (anonymous):

ok so it would be 120

OpenStudy (anonymous):

\[360+240=600,600\div 5=120\] right

OpenStudy (anonymous):

so 2nd answer is \[3(\cos (120°) + i \sin (120°))\]

OpenStudy (anonymous):

lather, rinse, repeat

OpenStudy (anonymous):

alright. so once again i add 360 and divide by 5?

OpenStudy (anonymous):

\[600+360=960, 960\div 5=192\]

OpenStudy (anonymous):

3(cos(192°)+isin(192°))

OpenStudy (anonymous):

third answer is \[3(\cos (192°) + i \sin (192°))\] you are too fast for me

OpenStudy (anonymous):

haha sorry

OpenStudy (anonymous):

no that is good, you get the picture, even though evidently you will never use this again

OpenStudy (anonymous):

i will let you finish the last two \[960+360\] etc

OpenStudy (anonymous):

ok thanks so much for the help. thats everything i needed help with

OpenStudy (anonymous):

you are welcome, hope you pass